A bullet is fired from a rifle that is held 1.5m above the ground in a horizontal position. The Initial speed of the bullet is 1210 m/s. find (a) the time it takes for the bullet to strike the ground (b) the horizontal distance traveled by the bullet
1
Expert's answer
2011-10-06T09:14:23-0400
(a) the time ittakes for the bullet to strike the ground
gt^2/2=1.5 => t=sqrt(3/g)=0.553 s
(b) the horizontal distance traveled by thebullet
S=1210*0.553=669.13 m